OpenStudy (anonymous):

Find the vertex of the parabola by using the vertex formula. f(x)=x^2+8x-9

OpenStudy (anonymous):

f(x) = x^2 - 8x - 9 By using completing the squares: f(x) = (x-4)^2 - 25 Vertex = (4, -25) As the coefficient of the x^2 is positive, it is a u-shaped quadratic curve.